solve (y-xz)p+(yz-x)q=(x+y)(x-y)


1
Expert's answer
2021-01-07T18:14:24-0500

Given, "(y-xz)p+(yz-x)q = (x+y)(x-y)".

"\Rightarrow(y-xz)p+(yz-x)q = x^2-y^2"


This equation is of the form "Pp+Qq=R" (Lagrange's linear partial differential equation).

Here, "P = y-xz, Q=yz-x, R = x^{2}-y^{2}".


The subsidiary equations are

"\dfrac{dx}{P}=\dfrac{dy}{Q}=\dfrac{dz}{R}\\
\dfrac{dx}{y-xz}=\dfrac{dy}{yz-x}=\dfrac{dz}{x^{2}-y^{2}}~~~~~~~~~~~-(1)".

Each ratio of (1) is equal to "\dfrac{xdx+ydy}{(y^{2}-x^{2})z} = \dfrac{dx+dy}{(1+z)(y-x)}".

Let us consider,

"\dfrac{xdx+ydy}{(y^{2}-x^{2})z} = \dfrac{dz}{x^{2}-y^{2}}\\\Rightarrow
xdx+ydy+zdz=0\\
\text{Integrating, }\\ \int xdx + \int ydy + \int zdz=0\\\Rightarrow
\dfrac{x^{2}}{2}+\dfrac{y^{2}}{2} +\dfrac{z^{2}}{2}=\dfrac{c_{1}}{2}\\\;\\\Rightarrow
x^{2}+y^{2}+z^{2}=c_{1}"


Consider,

"\dfrac{dx+dy}{(1+z)(y-x)}=\dfrac{dz}{(x+y)(x-y)}\\\Rightarrow
\dfrac{dx+dy}{1+z}=\dfrac{dz}{-(x+y)}\\\Rightarrow
-(x+y)d(x+y)=(1+z)dz\\
\text{Integrating,}\\
-\int (x+y)d(x+y)=\int (1+z)dz\\\Rightarrow
-\dfrac{(x+y)^2}{2}=\dfrac{(1+z)^{2}}{2}-\dfrac{c_{2}}{2}\\\;\\
\Rightarrow
(1+z)^{2}+(x+y)^{2}=c_{2}"


The general solution is,

"\phi(c_{1},c_{2})=0\\
\phi\left(x^{2}+y^{2}+z^{2},\left((1+z)^{2}+(x+y)^{2}\right)\right)=0."